OpenStudy (anonymous):

a projectile is launched with a speed of 140 m/s at an angle of 45 degrees above the horizontal, from a height of 1500 above the ground.how far does it travel in the horizontal direction before hitting the ground.

OpenStudy (anonymous):

there are different ways of solving this problem. the way would do: equation of motion on y direction: \[y=v_{0y}t-gt^2/2\] giving y=-1500m, you can find time when the object hits the ground, and then use \[x=v_{0x}t\] to find displacement in horizontal direction: \[-1500=140*\sqrt{2}/2t-9.81/2t^2\] move things around: \[4.9t^2-99t-1500=0\] this is second order equation in the form ax^2+bx+c=0 solution: \[t_{1,2}=(-b\pm\sqrt{b^2-4ac})/2a\] you should get minus for one of t's, that is not possible. Thus use other one, and put in \[x=v_xt\]

OpenStudy (anonymous):

what about the other approach?

OpenStudy (anonymous):

can we make use parabola equation since the graph of projectile is simply a parabola?

OpenStudy (anonymous):

what do you mean by using parabola equation?

OpenStudy (anonymous):

starting from back of posted solution, at the end you need to solve \[x=v_x*t\] here you can find t by different ways. second way would be: you can think this motion in 2 parts, first from starting point to maximum height it would go \[t_{tomax}=v_{0y}/g\] then from top it does free fall to ground. distance it goes up to max: \[h_{tomax}=v_{0y}^2/2g\] time of free fall from max to ground \[t_{from max to ground}=\sqrt{2(h_{tomax}+1500)/g}\] then \[x=v_{0x}(t_{tommax}+t_{from max to ground})\]

OpenStudy (anonymous):

sry what is √2/2t in the equation?

OpenStudy (anonymous):

are you asking: \[t=\sqrt{2h/g}\] it is from free fall equation: \[h=gt^2/2\]

OpenStudy (anonymous):

i mean in −1500=140∗√2/2t−9.81/2t^2, i know 140 is V0y but what about √2/2t?

OpenStudy (anonymous):

o you need \[v_0y=v_0\sin{\theta}\]

OpenStudy (anonymous):

\[\sin{45}=\sqrt{2}/2\]

OpenStudy (anonymous):

OH!!!! I C I C thx!!!